OpenStudy (anonymous):

f(r)=6r+9 f(5)= f(-3)= f(6.8)=

OpenStudy (anonymous):

substitute the values for r in each equation

OpenStudy (anonymous):

f(5)=6(5)+9=39

OpenStudy (callisto):

f(5)= 6(5)+9 =39 f(-3)=6(-3)+9 =-9 try the last one!

OpenStudy (anonymous):

f(5) = 39 f(-3)= -9 f(6.8) = 49.8

OpenStudy (anonymous):

f(6.8)=6(5)+9=39-39=0 ??

OpenStudy (callisto):

no! f(6.8)=6(6.8)+9=49.8

OpenStudy (anonymous):

oh, I see what I did... oops

OpenStudy (callisto):

you should replace 'r' with the number in the blanket and calculate it